OpenStudy (anonymous):

x/2-y/3=5/6 x/5-y/4=29/10

OpenStudy (texaschic101):

are these separate problems ?

OpenStudy (texaschic101):

when your dealing with fractions in an equation, and you don't want to, just multiply the entire equation by the common denominator. x/2 - y/3 = 5/6 --- common denominator is 6 (6)(x/2) - (6)(y/3) = (6)(5/6) 3x - 2y = 5 same with this one.. x/5 - y/4 = 29/10 -- common denominator is 20 (20)(x/5) - (20)(y/4) = (20)(29/10) 4x - 5y = 58 now if its one problem and your solving for x and y, you have to let me know and I can continue.